OpenStudy (anonymous):

What is the solution to the system of equations? 5x+4y=8 x-2y+10

OpenStudy (anonymous):

you have to find x and y.

OpenStudy (anonymous):

you have to equal x or y to solve so, 5x+4y=8 x-2y+10 (2) 5x+4y=8 2x-4y=20 ======= 7x=28 x=28/7 x=4 +4y and -4y will be cancelled.

OpenStudy (anonymous):

to find y,you use any one original equation and plug in the value of x 5(4)+4y=8 20+4y=8 4y=8-20 4y=-12 y=-12/4 y=-3

OpenStudy (anonymous):

@Riot_Lion01 Understand?

OpenStudy (anonymous):

uhhhhh i think so

OpenStudy (anonymous):

Okay.. :)

OpenStudy (anonymous):

so would y = -3 and x = 4 or do I still have to plug it in?

OpenStudy (anonymous):

to the equation

OpenStudy (anonymous):

uhh bye i guess

OpenStudy (anonymous):

Do you know @tianalovesyou17

OpenStudy (anonymous):

so 4,-3?

OpenStudy (anonymous):

Yeah that's the answer.

OpenStudy (anonymous):

x=4.y=-3

OpenStudy (anonymous):

Yes, the solution to the system is (4, -3). You can even plug the numbers in to one of the original equations to see for yourself.

OpenStudy (anonymous):

ok cool thank you :)

OpenStudy (anonymous):

@BunnyBree can you help me with a few more?

OpenStudy (anonymous):

I'll try. (: